Hello out there,I'm writing about some problems that I need help with and thanks to whoever can help,the first one is,Determine the slope and y-intercept of the graph :3x+y=8.The next one is,Determine the slope and y-intercept of the graph:-x+2y=4.The next one is,Write an equation in slope-intercept form of a line with the following characteristics:Perpendicular to the graph of:y=3/2x-3 with a y-intercept of 5.And the last one is,Write an equation in slope-intercept form of a line with the following characteristics:parallel to the graph of:-4x+3y=12 and passes through the origin.Thanks for any help that I can get.
Answer by stanbon(75887)

Determine the slope and y-intercept of the graph :3x+y=8.
Rewrite as y=-3x+8
slope = 3; y-intercept = 8
------------------
The next one is,Determine the slope and y-intercept of the graph:-x+2y=4.
Rewrite as y=(1/2)x+4
slope = 1/2
y-intercept = 4
----------------
The next one is,Write an equation in slope-intercept form of a line with the following characteristics:Perpendicular to the graph of:y=3/2x-3 with a y-intercept of 5.
The given line has slope = 3/2; the perpendicular line will have slope = -2/3.
EQUATION:
y=(-2/3)x+5
-----------------------
And the last one is,Write an equation in slope-intercept form of a line with the following characteristics:parallel to the graph of:-4x+3y=12 and passes through the origin.
Given line has slope 4/3
EQUATION: y=(4/3)x
====================
Cheers,
Stan H.